Find the slope of the line that is parallel ×-8y=5
Found 2 solutions by stanbon, Cromlix:
Answer by stanbon(75887)

Find the slope of the line that is parallel ×-8y=5
------
8y = x - 5
----
y = (1/8)x - (5/8)
-----
slope = 1/8
=================
Cheers,
Stan H.
================

Answer by Cromlix(4381)

Put in y = mx + c form
x - 8y = 5
-8y = - x + 5
Multiply through by -1
8y = x - 5
Divide through by 8
y = 1/8x - 5/8
As lines that are parallel have
the same gradient.
m1 = m2
So, slope of line that is
parallel to x - 8y = 5
= 1/8
Hope this helps.
:-)